QUANTUM SOLVATION OF CO_{2} BY He ATOMS: FROM ONSET TO NANODROPLET

\underline{HUI LI}, PIERRE-NICHOLAS ROY and ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

Spectroscopic studies of molecules embedded in helium droplets provide a unique opportunity to investigate a superfluid. High-resolution infrared spectra of CO_{2}-(He)_{N} clusters ranging in size from ‘small’ (N ≃1−17) to ‘intermediate’ (N ∼ 60) have been reported recently by McKellar and co-workers.^{a} However, they have not yet been able to extend this range to clusters large enough that the observed vibrational band origins shifts Δν_{3} and rotational constants B_{0} and D_{0} approached the nanodroplet limit values.^{b} This paper describes the use of modeling techniques to explore the domain between intermediate and ‘larger’ (N ∼10^{2} − 10^{3}) clusters, seeking to characterize the onset of nanodroplet behaviour. Path integral Monte Carlo (PIMC) simulations using the worm algorithm can in principle provide reliable quantitative results for “larger” clusters approaching the nanodroplet limit. We have recently determined a three-dimensional analytical ‘Morse/Long-Range’ potential energy surface for CO_{2}-He, which explicitly depends on the Q_{3} asymmetric-stretch vibrational motion of CO_{2}, and takes account of change in the average value of Q_{1} when ν_{3} is excited.^{c}^{,}^{d} Using this potential, predicted vibrational frequency shifts have been found to be in excellent agreement with experiment across the range N =1 − 40,^{d} so it is expected to provides a good description of larger clusters. This paper describes results obtained on using this potential in PIMC simulations to predict the effective rotational constant B_{0} and the Δν_{3} band-origin shift for CO_{2} in (He)_{N} clusters with N > 10^{2}. The importance of identical-particle exchange effects will be examined by comparing results obtained using both Boltzmann and Bose-Einstein statistics.

a J. Tang, A.R.W. McKellar, F. Mezzacapo and S. Moroni Phys. Rev. Lett. 92, 145503 (2004); A.R.W. McKellar, J. Chem.

Phys. 128, 044308 (2008).

b K. Nauta and R.E. Miller, J. Chem. Phys. 115, 10254 (2001); R. Lehnig and W. Jäger, Chem. Phys. Lett. 424, 146 (2006).

c H. Li and R.J. Le Roy, J. Chem. Chem. Phys. 10, 4128 (2008).

d H. Li, N. Blinov, P.-N. Roy and R.J. Le Roy, J. Chem. Phys. (2009, in press).